Solve



(D^(2)-DD'-2D'^(2)+2D+2D')z=sin(2x+y)

1
Expert's answer
2022-01-09T11:26:18-0500

"(D^2-DD'-2D'^2+2D+2D')z=sin(2x+y)"


auxillary equation:

"m^2-m-2+2m+2=0"

"m^2+m=0"

"m_1=0,m_2=-1"

complementary function:

"C.F.=f_1(y+m_1x)+f_2(y+m_2x)=f_1(y)+f_2(y-x)"


 particular integral:

"P.I.=\frac{1}{F(D,D')}sin(ax+by)=I.P.\frac{1}{F(D,D')}e^{i(ax+by)}=I.P.\frac{1}{F(ia,ib)}e^{i(ax+by)}"


"P.I.=I.P.\frac{1}{(2i)^2-2i\cdot i-2i^2+2\cdot2i+2i}e^{i(ax+by)}=I.P.\frac{1}{6i}(cos(2x+y)+isin(2x+y))="


"=I.P.-\frac{i}{6}(cos(2x+y)+isin(2x+y))=-\frac{1}{6}cos(2x+y)"


"z=C.F.+P.I.=f_1(y)+f_2(y-x)-\frac{1}{6}cos(2x+y)"